# CountZES: Counting via Zero-Shot Exemplar Selection

### v1

@@ 5 Conclusion
We presented CountZES, a training-free framework for ZOC, which formulates exemplar discovery as a three-stage process (DAE, DGE, and FCE), each contributing complementary cues from semantic, statistical, and visual perspectives. DAE ensures text-grounded semantic alignment, DGE introduces density-driven numerical consistency, and FCE enforces feature-level consensus to identify representative exemplar. This naturally induces diversity in the exemplar set, enhancing counting performance. Extensive experiments across natural, aerial, and medical benchmarks demonstrate its superior performance compared to SOTA. CountZES challenges the conventional reliance on task-specific training and supervision, posing an open question: Can principled exemplar selection alone be the key to scalable, domain-resilient zero-shot object counting?
Figure 12: Illustration of the SSES module within the DAE stage. Given the detector-derived coarse box, SSES leverages the text-conditioned CLIP similarity map to identify high-confidence response peaks using an adaptive percentile-based relaxation strategy. These peaks are used as point prompts to SAM, generating multiple candidate masks. Each mask is evaluated using a composite score that balances semantic alignment and spatial purity, and the highest-scoring mask is selected as the refined single-instance exemplar. The visual pipeline highlights how similarity-driven peak sampling and mask scoring together transform a noisy coarse region into a precise exemplar.
@@ A.1 Ablation on $k$ peaks in SSES module
Similarity-guided SAM-based Exemplar Selection (SSES) in Detection-Anchored Exemplar (DAE) stage, refines the initial-box selection into a clean, single-instance exemplar by jointly leveraging text-conditioned similarity map and SAM’s segmentation. It first locates the most text-aligned subregions within the coarse box by detecting high-confidence peaks in the similarity map using an adaptive percentile-relaxation strategy, ensuring robust peak detection under both strong and weak similarity patterns. As illustrated in Fig. 11 and 12, each detected peak serves as a point prompt to SAM, producing a set of candidate masks that are then evaluated using a composite score combining semantic alignment (mean percentile rank of similarities within the mask) and spatial purity (normalized entropy). This scoring mechanism prioritizes masks that are simultaneously text-consistent and instance-coherent, allowing SSES to reliably isolate single-object exemplars even in cluttered or ambiguous detections. We empirically set the maximum number of peaks at $k=16$, which provides the best trade-off between coverage and precision for counting performance, with an ablation over k=4,8,16k={4,8,16} reported in Table 6.
[…]
CountZES ($k$ = 4) 16.41 91.47
CountZES ($k$ = 8) 16.22 91.65
[…]

### v2

@@ 5 Conclusion
We presented CountZES, an inference-only framework for ZOC, which formulates exemplar discovery as a three-stage process (DAE, DGE, and FCE), each contributing complementary cues from semantic, statistical, and visual perspectives. DAE ensures text-grounded semantic alignment, DGE introduces density-driven numerical consistency, and FCE enforces feature-level consensus to identify representative exemplars. This induces diversity into the exemplar set, thereby enhancing counting performance. Extensive experiments across multiple domains demonstrate its generalization and superior performance compared to SOTA.
Broader Impact: This paper presents work whose goal is to advance the field of machine learning. At present, we do not anticipate any discernible adverse ethical or social implications of our work.
@@ A.1 Ablation on $k$ peaks in SSES module
Similarity-guided SAM-based Exemplar Selection (SSES) in Detection-Anchored Exemplar (DAE) stage, refines the initial-box selection into a clean, single-instance exemplar by jointly leveraging text-conditioned similarity map and SAM’s segmentation. It first locates the most text-aligned subregions within the coarse box by detecting high-confidence peaks in the similarity map using an adaptive percentile-relaxation strategy, ensuring robust peak detection under both strong and weak similarity patterns. As illustrated in Fig. 8 and 9, each detected peak serves as a point prompt to SAM, producing a set of candidate masks that are then evaluated using a composite score combining semantic alignment (mean percentile rank of similarities within the mask) and spatial purity (normalized entropy). This scoring mechanism prioritizes masks that are simultaneously text-consistent and instance-coherent, allowing SSES to reliably isolate single-object exemplars even in cluttered or ambiguous detections. We empirically set the maximum number of peaks at $k=16$, which provides the best trade-off between coverage and precision for counting performance, with an ablation over k=4,8,16k={4,8,16} reported in Table 6.
Figure 9: Illustration of the SSES module within the DAE stage. Given the detector-derived coarse box (in yellow), SSES leverages the text-conditioned CLIP similarity map to identify high-confidence response peaks using an adaptive percentile-based relaxation strategy. These peaks are used as point prompts to SAM, generating multiple candidate masks. Each mask is evaluated using a composite score that balances semantic alignment and spatial purity, and the highest-scoring mask is selected as the refined single-instance exemplar. The visual pipeline highlights how similarity-driven peak sampling and mask scoring together transform a noisy coarse region into a precise exemplar.
@@ A.2 P2P Prompting and Single-Instance Filtering
Method MAE RMSE
CountZES ($k$ = 4) 16.41 91.47
CountZES ($k$ = 8) 16.22 91.65
CountZES ($k$ = 16) 15.77 91.40
A.3 Mathematical Details
Entropy Computation. For each candidate detection box $\mathbf{b}_{i}=(x_{0},y_{0},x_{1},y_{1})$, we extract the CLIP similarity values from the similarity map $S\in[0,1]^{H\times W}$ within the box region, forming a set $\mathcal{V}_{i}=\{S(x,y)\mid(x,y)\in\mathbf{b}_{i}\}$. To estimate the distribution of similarity responses, we construct a histogram with $B=20$ equally spaced bins over the range $[0,1]$, yielding an empirical probability mass function $h_{i}(k)$ for $k=1,\dots,B$.
We compute the Shannon entropy of this distribution as
$H(\mathbf{b}_{i})=-\sum_{k=1}^{B}\tilde{h}_{i}(k)\log\tilde{h}_{i}(k),$ (8)
where $\tilde{h}_{i}(k)=\frac{h_{i}(k)+\epsilon}{\sum_{j=1}^{B}(h_{i}(j)+\epsilon)}$ applies Laplace smoothing with $\epsilon=10^{-8}$ for numerical stability. Intuitively, low entropy corresponds to spatially consistent and semantically coherent regions, while high entropy indicates mixed or noisy content. To enable fair comparison across candidate boxes, we further perform min–max normalization over all entropy values:
$\widehat{H}(\mathbf{b}_{i})=\frac{H(\mathbf{b}_{i})-H_{\min}}{H_{\max}-H_{\min}+\epsilon},$ (9)
where $H_{\min}$ and $H_{\max}$ denote the minimum and maximum entropy among valid detections, respectively.
Pseudo-GT Count Estimation. Let $\mathcal{B}_{\text{single}}=\{\mathbf{b}_{i}\}_{i=1}^{N}$ denote the set of candidate single-instance boxes obtained after single instance filtering via density-based ROI count. For each candidate box $\mathbf{b}_{i}$, we obtain a local count estimate by leveraging the pretrained off-the-shelf density estimator $f_{\text{cnt}}(\cdot)$:
$\hat{c}_{i}=f_{\text{cnt}}(I,\mathbf{b}_{i}),\quad i=1,\dots,N,$ (10)
where $\hat{c}_{i}\in\mathbb{Z}$ represents the predicted number of instances within $\mathbf{b}_{i}$. These per-box predictions are noisy due to imperfect localization and counter uncertainty.
To robustly aggregate these estimates, we model their empirical distribution using kernel density estimation (KDE), yielding a continuous probability density function:
$\hat{p}(c)=\frac{1}{Nh}\sum_{i=1}^{N}K\!\left(\frac{c-\hat{c}_{i}}{h}\right),$ (11)
where $K(\cdot)$ is a Gaussian kernel and $h$ denotes the bandwidth parameter. The pseudo ground-truth count is then defined as the mode of this distribution:
$c_{\text{pseudo}}=\arg\max_{c}\,\hat{p}(c).$ (12)
In cases where the distribution exhibits negligible variance or insufficient distinct values, we adopt a robust fallback by assigning $c_{\text{pseudo}}$ to the common value if all $\hat{c}_{i}$ are identical, or to the rounded mean $\lfloor\frac{1}{N}\sum_{i=1}^{N}\hat{c}_{i}\rceil$ otherwise. This aggregation strategy consolidates multiple uncertain predictions into a stable reference count while mitigating the influence of outliers.
Figure 11: Performance–runtime tradeoff among inference-only zero-shot counting methods, where MAE is plotted against average inference time (sec/image) and bubble size represents RMSE. CountZES achieves the best accuracy with competitive inference efficiency.
A.4 Datasets
FSC-147 dataset (Shi et al., 2022) contains 6,135 paired images across 147 object classes and is specifically designed for class-agnostic counting. It is split into 3,659 training, 1,286 validation, and 1,190 test images with non-overlapping classes across splits, making it well-suited for evaluating zero-shot object counting. CARPK (Hsieh et al., 2017) dataset provides a bird’s-eye view of 89,777 cars across 1,448 parking-lot images, serving as a standard benchmark for assessing cross-domain generalization from natural to aerial scenes. The recently introduced PerSense-D dataset(Siddiqui et al., 2025) serves as a dense counting benchmark, comprising 717 images spanning 28 object categories, with an average of 53 target instances per image and a total of 36,837 annotated objects across the dataset. VGG dataset (Lempitsky and Zisserman, 2010) includes 200 fluorescence-light microscopy images containing simulated bacterial cells, each image featuring approximately 174 ± 64 overlapping cells at varying focal depths to emulate realistic microscopic imaging conditions. Finally, MBM (modified bone marrow) dataset (Paul Cohen et al., 2017) contains 44 microscopic images with 126 ± 33 nuclei per image, representing challenging biological counting scenarios with high appearance variability.
A.5 Additional Qualitative Results
Fig. 12 presents additional qualitative results of CountZES on the FSC-147 dataset, showcasing its exemplar selection and counting performance across varied object densities and appearances.
Figure 12: Additional qualitative results of CountZES on FSC-147 dataset.
A.6 Runtime Efficiency
We analyze the runtime efficiency of CountZES on FSC147 dataset, with all experiments conducted on a machine equipped with an NVIDIA RTX A5000 GPU. Figure. 11 illustrates the tradeoff between inference efficiency and counting accuracy among inference-only zero-shot counting methods, where MAE is plotted against average inference time and bubble size denotes RMSE. SAM exhibits high error despite moderate runtime (3.4 sec/image), while methods such as TFCAC incur substantial computational overhead (5.26 sec/image) without achieving corresponding accuracy improvements. TFOC achieves faster inference (2.1 sec/image) but suffers from noticeably higher error. In contrast, CountZES occupies the optimal region of the tradeoff space, achieving the lowest MAE and RMSE with competitive inference time (2.3 sec/image). These results demonstrate that the proposed multi-stage exemplar selection pipeline yields substantial performance gains without incurring very high computational cost, effectively balancing accuracy and efficiency in an inference-only setting.
A.7 Limitations
While CountZES demonstrates better generalization across domains, certain limitations stem from the use of off-the-shelf pre-trained components like density estimators and SAM. In cross-domain settings, where the pretrained counter is trained on distributions different from the target domain, this can lead to noise in peak detection via P2P prompting. Also, as vanilla SAM lacks semantic awareness, a point prompt may yield only partial masks, capturing fragments of the object rather than complete instances. Although the scoring-based selection mechanism generally favors the most reliable exemplar at each stage, residual noise can still propagate when the majority of candidate exemplars are imperfect. As illustrated in Fig. 13 for the class “watches,” a point placed on the watch frequently segments only the dial while excluding the wristband, resulting in incomplete exemplar boxes. Additional examples in Fig. 14 highlight similar failures: for “nail polishes,” the mask captures only part of the bottle; for “comic books,” the prompt segments characters on the cover instead of the whole book; for “eggs,” SAM isolates only the yolk region; and for “sauce bottles,” it focuses on the label instead of the full object. Although the scoring mechanism generally prefers the best available candidates, residual noise can still propagate when most exemplars are imperfect. Nonetheless, a key strength of CountZES lies in its modular architecture. Each component, counter or segmentation backbone, can be seamlessly replaced with more advanced alternatives without altering the overall pipeline. This flexibility opens promising avenues for further enhancing cross-domain robustness and improving exemplar fidelity in future extensions.
Figure 14: Additional failure cases highlighting CountZES selection of incomplete boxes as exemplars. For “nail polishes,” SAM-based point prompt segmentation in CountZES captures only partial objects; for “comic books,” point prompts isolate characters or cover details instead of the full book; for “eggs,” SAM segments only the yolk region; and for “sauce bottles,” it segments only the label instead of the whole bottle. These examples show how imperfect masks can impair downstream exemplar selection in CountZES.